OpenStudy (awesomeadam65):

You Get a Medal If you answer this question PLZ HELP Which of these is a rational number A. 150 B. 441 C. 200 D. 250

OpenStudy (awesomeadam65):

@IQ250 @jonnnnnnnnxyyyyyyyyyy

OpenStudy (awesomeadam65):

HELP ME PLZ

OpenStudy (anonymous):

is this a trick question? are you sure you're not missing any symbols? they all look like rational numbers to me

OpenStudy (awesomeadam65):

here what they are suppose to look like

OpenStudy (nothingwasthesame):

only with b can you get a rational square root

OpenStudy (awesomeadam65):

\[\sqrt{150} \sqrt{441} \sqrt{200} \sqrt{250}\]

OpenStudy (nothingwasthesame):

you need to show work